# The role of spectator modes in the quantum-logic spectroscopy of single trapped molecular ions

### v1

@@ 4 Results and discussion
The simulations shown in Fig. 3 (a) and (b) indicate that upon applying the ODF, the average phonon number of the target mode, , increases first quadratically and then the gradient is getting reduced. The excitation rate depends on the ac-Stark shift which is directly proportional to the resultant ODF [14]. Additionally, the results of the simulations is compared to the outcome when the motional excitation is solely produced by a displacement operator with the coherent-state time-evolution defined as (dotted red trace in Fig. 3 (a)). Such a displacement model results from a first-order approximation of the Hamiltonian Eq. (9) and is often used to describe coherent motional excitation [17, 38]. However, while this treatment describes the early stages of the time evolution satisfactorily, it is not sufficient to describe the excitation for ODF pulse times 200 s and exceeding 15 phonons. This result underlines the importance of taking higher-order terms in the expansion of the Hamiltonian.
@@
Moreover, it is instructive to compare the differences between the classical and quantum simulations with no spectator mode population for the same ac-Stark shift of 16 kHz, i.e. dotted black and solid blue traces in Fig. 3 (a). While both treatments yield similar results at the beginning of the dynamics, the outcomes only slightly diverge with increasing interaction time. This finding indicates the minute presence of quantum effects in the excitation dynamics for the considered excitation times. It suggests that, for fully suppressed populations in the spectator modes, classical treatment would be sufficient.
@@
Modelling population distribution in the spectator mode necessitates using quantum treatment. The presence of the spectator mode population affects the motional excitation through the Debye-Waller effect, reducing the effective motional excitation of the ions in the target mode, as shown in Fig. 3 (b). Since there is no cross-coupling between the target and spectator modes, and the modulation frequency of the running lattice is far detuned from the spectator modes, there is no increase in the spectator mode population. Consequently, in subsequent simulations of the experimental data, the spectator modes were assumed to exhibit constant thermal populations.
The evolution of the target-mode populations over different lattice interaction times is presented in Fig. 3 (c,d). For the early stages of the time evolution when the spectator mode was cooled to its ground state (panel c), the population distribution of the target mode is close to Poissonian (solid lines) which would be expected for a coherent state of the corresponding average phonon number. However, at moderate interaction times, when steadily increases, the distributions become noticeably narrower than Poissonian, indicating a more intricate excitation mechanism. If, by contrast, the spectator mode is excited (panel d), highly asymmetric distributions are obtained.
@@
To make contact with experiment, Rabi oscillations measured on the blue sideband of the transition in Ca+ transition following an ODF pulse used to probe the internal state of the molecule are shown for different spectator-mode populations in Fig. 4. First, consider the effect of different populations in the radial modes while both axial modes were initially cooled close to their ground states (Fig. 4 (a)). It is evident that any effect of radial-mode populations is negligible at the sensitivity limit of the present experiments: no difference was observed whether both, only a single or none of the radial modes were ground-state cooled. This result is expected because the coupling (described by the Lamb-Dicke parameters) between the lattice laser beams propagating along the axial direction and the radial motional modes is negligible. If any influences of radial-mode populations would have been detected, they would have indicated other effects such as couplings between the axial and radial modes caused by, e.g., trap asymmetries.
@@
The results of the Rabi thermometry strongly depend on the populations of the axial modes, , as evidenced in Eq. (6). We carried out quantum simulations of the lattice interacting with N at different ac-Stark shifts and at the relevant experimental parameters (lattice time, temperature of the modes) from which were computed. The resulting theoretical Rabi flops were compared with the experimental data. The exact ac-Stark shift was extracted from the best match of the simulations when ground-state cooling both axial modes with the experiment.
@@
In Fig. 5, experimental Rabi flops (a) after initially cooling both axial modes close to their ground states and (b) no secondary cooling of the ax-OP mode following Doppler cooling (SB in Tab. 1) are compared with simulations. By comparing both experiments, it can be seen that excitations in the ax-OP spectator mode markedly reduce the contrast of the Rabi oscillations, which is well reproduced by the simulations. The reason can be traced back to the DW factors Eqs. (15) and (16) which are effective in the excitation by the optical lattice described by Eq. (3). The ac-Stark shift of kHz on N (from a single lattice beam) used in the simulations is close to the (17.5 1.0) kHz calculated for the lattice-beam intensities (2.30 0.12 W/mm2) estimated in the experiments [39].
@@
For the purpose of molecular-state identification, it is sufficient to probe the Rabi flop around the -time of the blue sideband where the highest signal-to-background ratio can be achieved [6]. A main motivation for improving cooling and understanding the excitation dynamics is to increase the contrast between signal and background (see, e.g., Fig. 2) and therefore improve the state-detection fidelities. The advantages are twofold. First, better cooling reduces the background signal. Second, the amplitude of the Rabi oscillations when cooling all normal modes close to their ground states was noticeably increased due to suppression of DW effects. For a Rabi -time of =20 s, the probability of detecting population transfer on the blue sideband with no lattice excitation (either background signal or N in excited state) was =0.05, while with the lattice excitation, when N was in the ground state, was =0.95. Following the treatment outlined in Ref. [6], a fidelity in state detection exceeding 99.99% is obtained under these conditions for as few as eight experimental repetitions. Thus, depending on the targeted fidelity, the number measurements and therefore the duty cycle of the experiment can be significantly reduced.
@@
The increase in sensitivity thus achieved opens avenues toward employing the present scheme in detecting multiple molecular states without the need to adjust the parameters of the optical lattice. This is illustrated in Fig. 6, where calculated ac-Stark shifts are shown for different rotational states of N and their individual hyperfine manifolds [31, 15]. Note that the REMPI scheme used to produce N is isomer-selective and produces only ions with nuclear spins quantum numbers and 2, and thus yields only even rotational states [21]. Hyperfine transitions from the rotational ground state, =0, for the laser frequencies used in this experiment, experience similar ac-Stark shifts (within 0.5 kHz) and are indicated by the blue trace. Almost all hyperfine transitions from =2, 4 exhibit a similar magnitude of the ac-Stark shift in the wavelength range shown and are indicated by the red-shaded area in Fig. 6. The ac-Stark shifts of hyperfine transitions from = 6 lie within the orange- and violet-shaded areas. The ac-Stark shifts for the states 0 are at least one order of magnitude weaker compared to =0 at the lattice wavelength indicated by the vertical black dashed line. Even higher excited rotational states, 6, are unlikely to be produced with the employed REMPI scheme.
@@
In order to detect motional excitation associated with population of these excited rotational states at the present lattice parameters which are optimised for detecting the rotational ground state, increasing the interaction time or the intensity of the lattice lasers is necessary. In Fig. 7, we report experiments which correspond to such a detection of excited rotational states. The ac-Stark shifts extracted from these measurements are compatible with the orange-shaded areas displayed in Fig. 6. The improved signal-to-background ratio after cooling all normal modes and an increased lattice-ion interaction time (here 2000 ) was a prerequisite for detecting the weak ODF produced by these higher rotational states at the present lattice parameters.
@@ 5 Conclusions
In conclusion, it was shown that spectator modes play an important role in the dynamics of motional excitation of a two-ion string in a trap with an optical lattice used for the quantum-state detection of molecular ions. Supported by simulations, we showed that this can be attributed to Debye-Waller effects in the interaction of the ions with the lattice. By cooling all motional modes of the two-ion string close to their ground states, we improved the signal-to-background ratio of the molecular state detection and achieved a fidelity exceeding 99.99% within only eight experimental repetitions. Also, the improved sensitivity of the current scheme opens up new possibilities for the identification of excited rotational states at lattice parameters optimised for the detection of the ground state.
@@
Acknowledgements We thank Mudit Sinhal and Gregor Hegi for their contributions to the development of the theory employed in this paper. The present work was supported by the Swiss National Science Foundation (grant nr. 200021_204123), the Eurpoean Partnership on Metrology (Funder ID: 10.13039/100019599, grant nr. 23FUN04 COMOMET), and the University of Basel.
@@
Consider the 1D model of a running optical lattice interacting with a two-ion string described by the Hamiltonian from Eq. (9). The motion of the ions can be described in terms of two normal modes – in-phase (denoted with ‘’) and out-of-phase (denoted with ‘’) mode, with the frequencies given in Eq. (3).
@@
Second, the dynamics of the two-ion crystal interacting with the modulated optical-dipole force is given by the Schrödinger equation:
@@
There there is no phonon exchange between the modes.
@@
with Lamb-Dicke parameters defined according to Eqs. (11-14). By rearranging this equation and using the definitions of DW factors from Eq. (5), we finally obtain Eq. (3). In the main text, the lattice excitation time is denoted with instead of .

### v2

@@ 4 Results and discussion
In Fig. 3 (a), the simulations for no spectator-mode population are shown. It can be seen that upon applying the ODF, the average phonon number of the target mode, , initially increases quadratically and then more slowly. In the early stage of the excitation when the ion is still in the Lamb-Dicke regime, , the excitation is equivalent to applying a displacement operator with the coherent-state time-evolution defined as , which is shown as a dotted red trace. Such a displacement model is a first-order approximation to the Hamiltonian from Eq. (9) and is often used to describe coherent motional excitation [16, 43]. However, while this treatment describes the early stages of the time evolution satisfactorily, when the motional state populations follow a Poissonian distribution, it is not sufficient to describe the excitation outside of the Lamb-Dicke regime. Then, it is important to consider higher-order terms in the expansion of the Hamiltonian, which are included in our simulations. Under the experimental conditions presented here, this becomes relevant already for ODF pulse times 200 s and 15 phonons. As the average motional quantum number increases, so does the spread of the wave packet. As a result, the motional excitation is reduced, and the motional-state distribution becomes narrower compared to a coherent state with the same average phonon population, as shown in Fig. 3 (c). Such narrowing of the phonon distribution for highly excited motional states is expected to yield high-contrast Rabi flops. Similar motional excitation outside the Lamb-Dicke regime was previously employed to create squeezed states of motion [44, 45].
@@
It is instructive to compare the differences between the classical and quantum simulations. Such a comparison with no spectator mode population for the same ac-Stark shift of 16 kHz is shown by the dotted black and solid blue traces in Fig. 3 (a). Both treatments yield similar results for the initial dynamics, and the outcomes only slightly diverge with increasing interaction time. This finding indicates the small role of quantum effects for the considered excitation times. The classical model correctly recognises the behaviour outside of the Lamb-Dicke regime, i.e. reduction of motional excitation, when the particle’s motion becomes comparable to the lattice wavelength. The agreement suggests that for fully suppressed populations in the spectator modes, a classical treatment would be sufficient.
From Figs. 3 (a) and (b), it is apparent that the excitation rate depends on the ac-Stark shift, which is directly proportional to the resultant ODF [13]. To detect smaller ac-Stark shifts, such as those originating from higher rotational states of N, it is required to increase the time of interaction with the lattice significantly to achieve comparable average final phonon populations.
By contrast, modelling the effect of population in spectator modes necessitates the use of a quantum treatment. The presence of the spectator-mode population affects the motional excitation through the Debye-Waller effect which reduces the effective motional excitation of the ions in the target mode, as shown in Fig. 3 (b). The motional-state distribution of the target mode then significantly deviates from a Poissonian, especially at higher motional excitations, as can be seen in Fig. 3 (d) for an ax-OP spectator mode population of = 8 phonons. Averaging over such a broad distribution of motional states significantly reduces the contrast of the resultant sideband Rabi flop that is important for reliable state detection.
@@
Note that since there is no cross-coupling between the target and spectator modes, and the modulation frequency of the travelling lattice is far detuned from the spectator modes, there is no increase in the spectator mode population. Consequently, in subsequent simulations of the experimental data, the spectator modes were assumed to exhibit constant thermal populations.
@@
To make contact with experiment, Rabi oscillations measured on the blue sideband of the transition in Ca+ transition following an ODF pulse used to probe the internal state of the molecule are shown for different spectator-mode populations in Fig. 4. First, consider the effect of different populations in the radial modes while both axial modes were initially cooled close to their ground states (Fig. 4 (a)). The minute change in Rabi frequency between the green and yellow/red traces can be attributed to the small but non-negligible difference in the ax-OP spectator mode population. In addition, the small loss of contrast around = 24 s may be related to differences in the final motional-state-population distribution between the cooling methods employed in the experiments (EIT cooling – yellow/red traces, SB cooling – green trace data). Besides these small differences, it is evident that any effect of radial-mode populations is negligible at the sensitivity limit of the present experiments: no major difference was observed whether both, only a single or none of the radial modes were cooled close to the ground-state. This result is expected because the coupling (described by the Lamb-Dicke parameters) between the lattice laser beams propagating along the axial direction and the radial motional modes is negligible. If any influences of radial-mode populations would have been detected, they would have indicated other effects such as couplings between the axial and radial modes caused by, e.g., the Kerr-like effects or trap asymmetries.
@@
The results of the Rabi thermometry strongly depend on the populations of the axial modes, , as evidenced in Eq. (6). We carried out quantum simulations of the lattice interacting with N at different ac-Stark shifts and at the relevant experimental parameters (lattice time, temperature of the modes) from which were computed. The resulting theoretical Rabi flops were compared with the experimental data. In Fig. 5, experimental Rabi flops (a) after initially cooling both axial modes close to their ground states and (b) no secondary cooling of the ax-OP mode following Doppler cooling (SB in Tab. 1) are compared with simulations. By comparing both experiments, it can be seen that excitations in the ax-OP spectator mode markedly reduce the contrast of the Rabi oscillations, which is well reproduced by the simulations. The reason can be traced back to the DW factors Eqs. (15) and (16) which are effective in the excitation by the optical lattice described by Eq. (3).
The contrast of the resulting Rabi flop, and therefore the state-detection fidelity, depends on the shape of the generated motional wave packet. Fig. 2 (c) shows the wave packets corresponding to the flops in Figs. 2 (a) and 5 (a). The dependence of the Rabi frequencies on the motional quantum number of the target mode is plotted in Fig. 2 (b). The high-contrast flop in Fig. 2 (a) originates from a narrow wave packet with an average phonon number in a region where the Rabi frequency is only weakly dependent on . Thus, the Rabi frequencies for excitations out of all states populated in the wave packet are similar which, by virtue of Eq. 6, results in a Rabi flop that nearly resembles an excitation out of a single state. Conversely, the flop in Fig. 5 (a) corresponds to a broader wave packet in Fock space averaging at where the dependence of the Rabi frequency on is more pronounced, leading to loss of contrast in the Rabi flop.
The exact ac-Stark shift in the simulation was extracted from the best match of the simulations with the experiment when both axial modes were cooled to their ground states. The ac-Stark shift of kHz on N (from a single lattice beam) used in the simulations is close to the (17.5 1.0) kHz calculated for the lattice-beam intensities (2.14 0.12 W/mm2) estimated in the experiments [46].
@@
For the purpose of molecular-state identification, it is sufficient to probe the Rabi flop around the -time of the blue sideband where the highest signal-to-background ratio can be achieved (interval indicated by the black dotted lines in Fig. 2 (a)) [5]. A main motivation for improving cooling and understanding the excitation dynamics is to increase the contrast between signal and background (see, e.g., Fig. 2 (a)) and therefore improve the state-detection fidelities. The advantages are twofold. First, better cooling of the axial modes reduces the background signal. Second, the amplitude of the Rabi oscillations when cooling normal modes close to their ground states was noticeably increased due to suppression of DW effects. The resulting state detection fidelity can be calculated using the experimental data from Fig. 2 (a) by following the treatment outlined in Refs. [13] and [5]. For a Rabi -time of =15.3 s, the probability of detecting population transfer on the blue sideband with no lattice excitation (either background signal or N in excited state) was =0.05, while with the lattice excitation, when N was in the ground state, was =0.95. Under these conditions, from the statistical analysis, we compute a state-detection fidelity exceeding 99.99% for as few as nine experimental repetitions. Note, however, that the actual state-detection fidelity achieved in the experiment may be lower because of systematic effects such as laser drifts and collisions with background gas (see also Ref. [47]). Thus, depending on the targeted fidelity, the number of measurements and therefore the duration of the experiment can be significantly reduced, e.g., comparing to Ref. [5], the experimental cycle improved by a factor of two. This is particularly important for room-temperature experiments with (molecular) species whose chemical lifetime is limited by reactions with background gases. Shorter exposure to the lattice lasers is also desirable due to the potential for off-resonant scattering of the ground molecular state. Both of these factors play a role in the experiments presented here.
@@
The increase in sensitivity thus achieved allows for detection of much smaller ac-Stark shifts which could originate from the population of rotationally excited states of N. In Fig. 6, the calculated ac-Stark shifts are shown for different rotational states of N and their individual hyperfine manifolds [46, 34, 14]. Note that the REMPI scheme used to produce N is isomer-selective and produces only ions with nuclear spins quantum numbers and 2, and thus yields only even rotational states [25]. Hyperfine transitions from the rotational ground state, =0, for the laser frequency used in this experiment, experience similar ac-Stark shifts (within 0.5 kHz) and are indicated by the blue trace. Almost all hyperfine transitions from =2, 4 exhibit ac-Stark shifts of a similar magnitude in the wavelength range shown and are indicated by the red-shaded area in Fig. 6. The ac-Stark shifts of hyperfine transitions from = 6 lie within the red- and violet-shaded areas. The ac-Stark shifts for the states 0 are at least one order of magnitude weaker compared to =0 at the lattice wavelength indicated by the vertical black dashed line. Higher excited rotational states with 6 are unlikely to be produced with the employed REMPI scheme. As demonstrated in Ref. [14], the range of possible states can be narrowed down by comparing the measured ac-Stark shifts with theory and taking advantage of the phase sensitivity of the effective ODF. In principle, the sensitivity for the detection of molecular states corresponding to smaller ac-Stark shifts can be further enhanced by increasing the lattice interaction time without the need to adjust other parameters of the optical lattice, such as it’s wavelength, intensity or polarization.
@@
In Fig. 7, we report experiments which correspond to such a detection of excited rotational states. The ac-Stark shifts extracted from these measurements are compatible with the orange-shaded area displayed in Fig. 6. We emphasise that the improved signal-to-background ratio after cooling all normal modes and an increased lattice-ion interaction time (here 2000 ) was a prerequisite for detecting the weak ODF produced by these higher rotational states at the present lattice parameters.
@@ 5 Conclusions
In conclusion, it was shown that spectator modes play an important role in the dynamics of motional excitation of a two-ion string in a trap with an optical lattice used for the quantum-state detection of molecular ions. Supported by simulations, we showed that this can be attributed to Debye-Waller effects in the interaction of the ions with the lattice. By cooling the spectator modes of the two-ion string close to their ground states, we improved the signal-to-background ratio in molecular-state detection and achieved a statistical fidelity exceeding 99.99% within only nine experimental repetitions. As a result, compared to our previous report [5], the total experimental time was reduced by a factor of two, in addition to the enhanced detection fidelity. This reduction is particularly valuable when probing short-lived molecular states and for minimizing off-resonant scattering from the lattice lasers. Furthermore, the improved sensitivity of the current scheme opens up new possibilities for the identification of excited rotational states at lattice parameters optimised for the detection of the ground state.
@@
Acknowledgements We thank Mudit Sinhal and Gregor Hegi for their contributions to the development of the theory employed in this paper. We also thank the referee for his insights that resulted in the addition of Figure 2 (b) and (c). The present work was supported by the Swiss National Science Foundation (grant nr. 200021_204123), the Eurpoean Partnership on Metrology (Funder ID: 10.13039/100019599, grant nr. 23FUN04 COMOMET), and the University of Basel.
@@
Data availability The primary data underlying the findings of the present study are available on Zenodo at https://zenodo.org/records/15234656.
Code availability The software codes used to generate the simulations in the present study are available on Zenodo at https://zenodo.org/records/15234656.
@@
Consider the 1D model of a travelling optical lattice interacting with a two-ion string described by the Hamiltonian from Eq. (9). The motion of the ions can be described in terms of two normal modes – in-phase (denoted with ‘’) and out-of-phase (denoted with ‘’) mode, with the frequencies given in Eq. (3).
@@
The dynamics of the two-ion crystal interacting with the modulated optical-dipole force is given by the Schrödinger equation:
@@
There is no phonon exchange between the modes.
@@
with Lamb-Dicke parameters defined according to Eqs. (11-14). By rearranging this equation and using the definitions of DW factors from Eq. (5), we finally obtain Eq. (3).